# exampleloeList of Examples

### v1

@@ 5.2.3. NLI
We use the 400M FAIR mNLI [^4] https://huggingface.co/facebook/bart-large-mnli model to predicts logits for entailment, contradiction, and neutral, and divide these logits by an MNLI temperature $T\in\{1,10,100\}$ As per the FAIR guidelines, we pass the temperature-scaled entailment and contradiction scores through a softmax layer and take the entailment probabilities.
@@ 5.3. Datasets
We obtain item descriptions from three real-world datasets: MovieLens25M [^5] https://grouplens.org/datasets/movielens/25m/, Yelp [^6] https://www.yelp.com/dataset, and Recipe-MPR (Zhang et al., 2023) (example item descriptions from each shown in Table 1 in the Supplementary Materials). After the filtering steps below for Yelp and MovieLens, we randomly sample 100 items to create $\mathbf{x}$. For Yelp, we filter restaurant descriptions to be from a single major North American city and to have at least 50 reviews and five or more category labels. For MovieLens, [^7] For all experiments with MovieLens, we use the 16k version of GPT-3.5-turbo-0613, due to MonoLLM requiring extra context length for $\mathbf{x}.$ we filter movies to be in the 10% by rating count with at least 20 tags, and let movie descriptions use the title, genre labels, and 20 most common user-assigned tags.
@@
Figure 8. The effect of user response noise on MAP@10 – error bars are 95% confidence intervals. At all noise levels, MonoLLM is outperformed by all PEBOL-P methods on Yelp and MovieLens and by UCB, Greedy, and ER on Recipe-MPR.
@@ 5.4. Research Questions
  • RQ1: How does PEBOL perform against the GPT-3.5 Mono-LLM baseline?
@@ 6. Experimental Results
Figure 4 shows MAP@10 over 10 dialogue turns for MonoLLM and PEBOL (UCB,TS,ER), with all methods starting near random guessing (reflecting a cold start) and showing clear preference learning over time. [^8] For each policy, we use the MNLI temperature that performed best on the most datasets (see Supplementary Materials) with continous responses. PEBOL achieves a 131% improvement over MonoLLM in MAP@10 after 10 dialogue turns on Yelp, an 88% improvement on MovieLens, and a 55% improvement on RecipeMPR. MonoLLM also displays a large drop in MAP@10 between turns 4 and 5 on Recipe-MPR which it takes an additional five turns to recover from. In contrast, we do not see any comparably large drops in our methods, indicating that the incremental Bayesian belief updating mechanism of PEBOL may make it less prone to catastrophic mistakes.
@@
Figure 6 compares PEBOL performance for PEBOL-B versus PEBOL-P. In general, PEBOL-P performed better, likely due to PEBOL-B discarding valuable information from entailment probabilities. Specifically Figure 6 compares PEBOL-B with each policy against PEBOL-P with the best performing policy on each dataset. On MovieLens, we see at least a 34% improvement in MAP from probabilistic entailment over binary entailment starting from turn 2, though we note that TS with binary updates performs comparably to the best probabilistic methods on Yelp.
To identify our best continuous method, we first compare the performance of PEBOL with PEBOL-P using MNLI temperatures 1, 10, and 100 with each policy. We keep the temperature with the highest MAP@10 at turn 10 for the most datasets: for Yelp and MovieLens, this is TS with MNLI temperature 1; for Recipe-MPR, it is UCB with MNLI temperature 10.
@@ 6.3. RQ3 - Effect of User Response Noise
Figure 8 shows the impact of user response noise on MAP@10 at turn 10 – PEBOL generally continues to outperform MonoLLM under user response noise. At turn 10 for all noise levels, MonoLLM is the worst performing method on Yelp and MovieLens, and on RecipeMPR it is always outperformed by UCB, Greedy, and ER. Methods typically maintain their relative performance under various noise levels – though an exception is MovieLens where TS is the best method under zero noise and nearly the worst one at 25% noise (though the margin between the best and worst method in this case is small).
@@ 6.5. RQ5 - Effect of Aspect History in Query Generation
We see an improvement in performance from including a list of previously generated aspects in the aspect generation prompt. In Figure 7, the average improvement across all turns and all datasets for including the aspect history is 50% for TS and 49% for UCB. Including the aspect generation history helps to avoid repeat queries, which gain no information and would frustrate a user.

### v2

@@ 5.2.4. NLI
We use the 400M FAIR mNLI [^6] https://huggingface.co/facebook/bart-large-mnli model to predicts logits for entailment, contradiction, and neutral, and divide these logits by an MNLI temperature $T\in\{1,10,100\}$ As per the FAIR guidelines, we pass the temperature-scaled entailment and contradiction scores through a softmax layer and take the entailment probabilities. We report PEBOL results using the best MNLI temperature for the most datasets.
@@ 5.3. Datasets
We obtain item descriptions from three real-world datasets: MovieLens25M [^7] https://grouplens.org/datasets/movielens/25m/, Yelp [^8] https://www.yelp.com/dataset, and Recipe-MPR (Zhang et al., 2023) (example item descriptions from each shown in Table 1 in the Supplementary Materials). After the filtering steps below for Yelp and MovieLens, we randomly sample 100 items to create $\mathbf{x}$. For Yelp, we filter restaurant descriptions to be from a single major North American city (Philadelphia) and to have at least 50 reviews and five or more category labels. For MovieLens, [^9] For all experiments with MovieLens, we use the 16k version of GPT-3.5-turbo-0613, due to MonoLLM requiring extra context length for $\mathbf{x}.$ we filter movies to be in the 10% by rating count with at least 20 tags, and let movie descriptions use the title, genre labels, and 20 most common user-assigned tags.
@@
Figure 8. The effect of user response noise on MRR@10 – error bars are 95% confidence intervals.
@@ 5.4. Research Questions
  • RQ1: How does PEBOL perform against the MonoLLM baselines?
@@ 6. Experimental Results
Figure 4 shows MRR@10 over 10 dialogue turns for MonoLLM and PEBOL (UCB,TS,ER), [^10] For each PEBOL policy, we use the MNLI temperature that performed best on the most datasets with continuous responses (see Supplementary Materials). with 95% confidence intervals (CIs) at turn 10 shown in Figure 8 (see Supplementary Materials for CIs for all turns and experiments). All methods start near random guessing, reflecting a cold start, and show clear preference learning over time.
Compared to GPT-MonoLLM, which uses the same LLM (GPT-3.5-turbo-0613) as PEBOL does for query generation, after 10 turns of dialogue PEBOL achieves: a mean MRR@10 of 0.27 vs. GPT-MonoLLM’s MRR@10 of 0.12 on Yelp; 0.18 vs 0.09 on MovieLens; and 0.17 vs 0.11 on Recipe-MPR, respectively. Compared to Gemini-MonoLLM, which uses a newer generation LLM (Gemini-Pro) than PEBOL for query generation, after 10 turns PEBOL still achieves a higher mean MRR@10 of 0.27 vs. Gemini-MonoLLM’s mean MRR@10 of 0.17 on Yelp; 0.18 vs. 0.15 on MovieLens, and 0.17 vs 0.16 on Recipe-MPR, respectively. While we did not have the resources to test PEBOL with Gemini query generation, we hypothesize that using a newer LLM for query generation can further improve PEBOL performance, since using the newer LLM (Gemini) shows performance improvements for the MonoLLM baseline.
@@
Figure 6 compares PEBOL performance using binary (PEBOL-B) vs. continuous (PEBOL-P) feedback, and shows that performance is typically better when continuous responses are used – indicating that binary feedback models discard valuable information from the entailment probabilities.
@@ 6.3. RQ3 - Effect of User Response Noise
Figure 8 shows the impact of user response noise on MRR@10 at turn 10 – PEBOL generally continues to outperform MonoLLM under user response noise. Specifically, at all noise levels, both MonoLLM baselines are outperformed by all PEBOL-P variants on Yelp, and by at least one PEBOL-P variant on MovieLens and Recipe-MPR.
@@ 6.5. RQ5 - Effect of Aspect History in Query Generation
As shown in Figure 7, we see improvements in PEBOL performance from including a list of previously generated aspects in the aspect generation prompt. For instance, the differences in mean MRR@10 from including vs. excluding the query history for TS after 10 turns were: 0.27 vs 0.16 for Yelp; 0.18 vs 0.14 for MovieLens, and 0.13 vs 0.09 for Recipe-MPR, respectively. Practically, including the aspect generation history also helps to avoid repeat queries, which gain no information and could frustrate a user.
@@ 7. Conclusion and Future Work
This paper presents a novel Bayesian optimization formalization of natural language (NL) preference elicitation (PE) over arbitrary NL item descriptions, as well as introducing and evaluating PEBOL, an algorithm for NL Preference Elicitation with Bayesian Optimization augmented LLMs. As discussed below, our study also presents many opportunities for future work, including for addressing some of the limitations of PEBOL and our experiment setup.
User Studies
Firstly, our experiments limited by their reliance on LLM-simulated users. While the dialogue simulations indicate reasonable behaviour in the observed results, such as initial recommendation performance near random guessing, preference learning over time, and coherent user responses in logs such as those shown in Figure 7, future work would benefit from human user studies.
Multi-Item Belief Updates
While the assumption that item utilities can be updated independently allows the use of a simple and interpertable Beta-Bernouilli update model for each item, it also requires a separate NLI calculation to be performed for each item, which is computationally expensive. A key future direction is thus to explore alternative belief state forms which enable the joint updating of beliefs over all items from a single NLI computation.
Collaborative Belief Updates
Since PEBOL does not leverage any historical interactions with other users, an important future direction is to study NL-PE which leverages collaborative, multi-user data. One possibility is to initialize a cold start user’s prior beliefs based on interaction histories with other users. Another direction is adapting collaborative filtering based belief updating, such as the methods used in item-based feedback PE techniques (e.g., (Christakopoulou et al., 2016)), to NL-PE.
Diverse Query Forms
While PEBOL uses a pointwise query generation strategy that selects one item description at a time for LLM context, future work can explore LLM-based acquisition functions with pairwise and setwise context selection. Such multi-item context selection would enable contrastive query generation that could better discriminate between item preferences.
NL-PE in ConvRec Architectures